40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 19 09:36:32 2013, Koji, Update, Green Locking, Xend green aligned ALS_OUTOFLOOP_130816.pdf
    Reply  Mon Aug 19 14:22:36 2013, rana, Update, Green Locking, Xend green aligned 
       Reply  Mon Aug 19 17:40:32 2013, Steve, Update, Green Locking, Xend green layout corrections beforeC.jpgnowC.jpgstillMultiple.jpg
          Reply  Mon Aug 19 19:08:35 2013, Koji, Update, Green Locking, Xend green layout corrections 
          Reply  Tue Aug 20 10:59:15 2013, Steve, Update, Green Locking, Xend green layout corrections faraway.jpgcloser.jpg
             Reply  Wed Aug 28 11:15:02 2013, Steve, Update, Green Locking, Xend green layout corrections 
    Reply  Mon Aug 19 16:18:56 2013, manasa, Update, Green Locking, Xend green aligned 
Message ID: 9025     Entry time: Mon Aug 19 09:36:32 2013     Reply to this: 9031   9033
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: Xend green aligned 

[Rana Koji]

This is an elog about the activity on Friday night.

- The X arm green beam was aligned with assist of the ASX system.

- M1 PZT alignment was swept while M2 PZT was under the control of ASX.

- Everytime M1 was touched, M2 was restored by manual alignment so that the REFL beam hits the center of the REFL PD.
  This way we could recover the lock of TEM00. Once TEM00 is recovered, ASX took care of the alignment of M2

- The error signal used by the cavity dither did not give us a good indication where the optimal alignment is.

- Thus the best alignment of M1 had to be manually scanned. The resulting maximum green transmission was ~0.88

- Once the beam was aligned, the out-of-loop stability of the Xarm was measured.
  There has been no indication of the improvement compared to Manasa's measurement taken before our beam alignment.

Attachment 1: ALS_OUTOFLOOP_130816.pdf  54 kB  | Hide | Hide all
ALS_OUTOFLOOP_130816.pdf
ELOG V3.1.3-